in 4 years John will be four thirds Graces age 5 years ago John was one year more than twice Grace's age how old are John Grace now?
Answer by hamsanash1981@gmail.com(151)

let Grace's present age be x, then after four years Grace age will be x+ 4
John's age will be = 4/3(x+4) and present age of John = 4/3(x+4)-4
five years ago Grace age will be x-5
John's age = present age -5 => 4/3(x+4)-4-5 => 4/3(x+4)-9 --- 1
Given that John's 5 years ago was one year more than twice the Grace's age
the John's age = 2(x-5)+1 ---2
Equating equation 1 and 2 we have
4/3(x+4)-9 = 2(x-5)+1
4/3(x+4) = 2x - 10 + 1 + 9
(4x + 16)/3 = 2x
4x + 16 =6x
16 = 2x
x = 8
Then john's persent age = 4/3(x+4)-4
4/3(12)-4
=12
therefore John and Grace are 12 and 8 years old